Show that dp/dt =_dv/dx
1
Expert's answer
2021-05-20T10:07:55-0400

The momentum of an object is "p=mv"

"\frac{dp}{dt}=\frac{d(mv)}{dt}"

"=m\frac{d(v)}{dt}"

"=m\frac{dv}{dt}(\frac{dx)}{dx})"

"=m(\frac{dx}{dt})(\frac{dv}{dx})
=mv(\frac{dv}{dx})" "since (\frac{dx}{dt})=v"

Therefore,

"\frac{dp}{dt}=mv(\frac{dv}{dx})"